3. (-/1.25 Points] DETAILS MY NOTES ASK YOUR TEACHER PRACTICE ANOTHER (a) What is the equivalent capacitance (in pF) of the gIf both questions can be answered I'd appreciate it thank you!
Answer #1

fon 4 +1 Ceq (parallel)=(1+2 Cea (series) = GC2 3 Citch tut 2 11 6 4 At 77 6 36/13 66147 t 66 (eg = 36 64 = 6.65 uf 13 cl (94. ने 3) टले ज 12 । 33.? । Ceq = 31 33-33 * 2 3) 33-33 2 10.62F